# Active Gaze Behavior Boosts Self-Supervised Object Learning

### v1

@@ Abstract
Due to significant variations in the projection of the same object from different viewpoints, machine learning algorithms struggle to recognize the same object across various perspectives. In contrast, toddlers quickly learn to recognize objects from different viewpoints with almost no supervision. Recent works argue that toddlers develop this ability by mapping close-in-time visual inputs to similar representations while interacting with objects. High acuity vision is only available in the central visual field, which may explain why toddlers (much like adults) constantly move their gaze around during such interactions. It is unclear whether/how much toddlers curate their visual experience through these eye movements to support learning object representations. In this work, we explore whether a bio-inspired visual learning model can harness toddlers’ gaze behavior during a play session to develop view-invariant object recognition. Exploiting head-mounted eye tracking during dyadic play, we simulate toddlers’ central visual field experience by cropping image regions centered on the gaze location. This visual stream feeds a time-based self-supervised learning algorithm. Our experiments demonstrate that toddlers’ gaze strategy supports the learning of invariant object representations. Our analysis also reveals that the limited size of the central visual field where acuity is high is crucial for this. We further find that toddlers’ visual experience elicits more robust representations compared to adults’ mostly because toddlers look at objects they hold themselves for longer bouts. Overall, our work reveals how toddlers’ gaze behavior supports self-supervised learning of view-invariant object recognition.
@@ 1 Introduction
Toddlers learn visual representations that support the recognition of object instances observed from different viewpoints within their first year of life (Kraebel & Gerhardstein, 2006; Ayzenberg & Behrmann, 2024). This early emergence of view-invariant recognition and the ease with which adults perform this skill hide the complexities of learning it. Images reaching the retina vary drastically when objects are turned in depth. Even state-of-the-art machine learning methods still make absurd recognition mistakes when faced with unusual viewpoints of objects (Dong et al., 2022; Abbas & Deny, 2023; Ruan et al., 2023). It raises the question of what learning mechanisms support such a view-invariant recognition in humans.
@@
One of the main theories of object recognition posits that the development of view-invariant object recognition rests on the brain’s ability to construct visual representations that slowly change over time (Földiák, 1991; Li & DiCarlo, 2008; Miyashita, 1988). The main idea is that learners abundantly manipulate (or walk around) objects while watching them, giving access to different views of a single object over a short period of time. By learning slowly changing representations, a brain discards rapidly changing information from an image (here, information about the view) and naturally builds view-invariant representations. Following this idea, recent computational studies proposed to simulate humans’ visual experience by generating or curating large-scale temporal sequences of rotating objects (Aubret et al., 2022; Schneider et al., 2021; Yu et al., 2023); they confirm that learning slowly changing representation may indirectly induce view-invariant object recognition. However, it is currently unclear if a toddler’s actual gaze behavior supports this learning mechanism.
@@
The significance of active gaze behavior stems from the limited area of high-acuity vision in humans. This area, known as the central visual field, covers only a few degrees of visual angle. This high-acuity processing is predominant in extracting semantic information in cortex areas related to object recognition (Quaia & Krauzlis, 2024; Yu et al., 2015). Such a small area of the visual field may be semantically unstable over time, as humans make three saccades per second on average. However, toddlers curate their own visual experience; compared to adults, objects held by toddlers appear bigger in the field of view due to their shorter arms (Bambach et al., 2018), their visual stimuli are simpler (Anderson et al., 2024) and their visual inputs semantically change on a slower timescale (Sheybani et al., 2023). The latter point may be critical to make slowness-based learning operational.
@@
In this paper, we explore whether a bio-inspired model of visual learning can utilize the actual eye-tracking derived visual experience of toddlers to develop robust object representations. For this, we leverage a dataset of head-camera recordings and gaze tracking from toddlers and adults during play sessions (Bambach et al., 2018). To simulate central visual experience, we crop image patches centered on tracked gaze locations. Then, we train a previously introduced self-supervised learning (SSL) model of temporal slowness (Schneider et al., 2021).
Our experiments demonstrate that toddlers’ gaze behavior supports the learning of view-invariant object representations within a single unsupervised 12-minute play session. Our analysis shows that: 1) toddlers’ gaze strategy boosts visual learning in comparison to several baselines; 2) modeling the central visual field crucially enables object learning. Furthermore, we discover that representations learned from toddlers’ visual experiences are also better than adults’, presumably because toddlers look longer at objects that they hold. In sum, our work provides a plausible model of how toddlers may efficiently learn view-invariant object recognition without supervision by exploiting their active gaze behavior during interactions with objects.
@@ 2 Related Work
Early computational studies found that slowness-based learning can extract representations of simple patterns that are invariant to position, size and rotation (Földiák, 1991; Wiskott & Sejnowski, 2002). Other works applied this principle to learn view-invariant object recognition (Wallis & Baddeley, 1997; Franzius et al., 2011; Einhäuser et al., 2005; Stringer et al., 2006). Recent advances in SSL allowed to scale the principle of temporal slowness to large sets of uncurated images of objects (Parthasarathy et al., 2022; Aubret et al., 2022; Schneider et al., 2021). This method was called SSL through time (SSLTT) (Aubret et al., 2022). On the machine learning side, SSLTT can boost category recognition (Aubret et al., 2024b; 2022; Sanyal et al., 2023), view-invariant object instance recognition (Schneider et al., 2021) and the alignment with human representations (Parthasarathy et al., 2023). On the bio-modeling side, SSLTT can shape human-like inter-object semantic similarities (Aubret et al., 2024a) and combines well with visuo-language SSL to model object learning during dyadic play (Schaumlöffel et al., 2023). However, all these approaches use curated, synthetic, or third-person data, leaving unclear whether the statistical structure of toddlers’ visual experience, combined with temporal slowness, can indeed support object recognition. Another notable work studied the learning of view-invariant object representations in impoverished visual environments through the eyes of reared chicken (Pandey et al., 2024). In contrast, we apply SSLTT on complex visual inputs extracted from head cameras carried by toddlers and/or adults during play sessions.
@@
There is a recent surge in trained machine learning models on egocentric video datasets, including models of temporal slowness. For instance, the large-scale Ego4d dataset (Grauman et al., 2022) has been used for training vision models (Nair et al., 2022; Ma et al., ; Anderson et al., 2022). However, egocentric videos for toddlers were missing (Anderson et al., 2022); this is a problem since existing research has found that the specific statistical structure of toddlers’ visual experience especially supports visual learning (Sheybani et al., 2024; Bambach et al., 2017; Sheybani et al., 2023). The SAYcam dataset presents longitudinal recordings of 150 hours (on average) from each of the three children participants (Sullivan et al., 2021). With SAYcam, computational studies have shown that SSL methods can learn category recognition, with/without temporal slowness (Orhan et al., 2020; Orhan & Lake, 2024; Orhan et al., 2024). Another related work studies whether the temporal and developmental structure of toddlers’ visual experience supports category and action recognition, through temporal slowness (Sheybani et al., 2024). Yet, these computational studies neglect the gaze location and its associated behavioral strategy, as their datasets do include the precise location of the individual’s gaze. We show in Section 4.2 that this is critical for learning good object representations.
@@
Our work extends previous approaches that also leverage the gaze location of a human to train vision models (Bambach et al., 2016; 2018). They also compare the quality of representations trained with toddlers’ versus adults’ experiences. However, these studies model the learning process with supervised learning, which is biologically implausible. This is important as, unlike bio-inspired self-supervised models that learn slowly changing representations, they are agnostic to the temporal structure of the visual experience, e.g., if toddlers look at an object for a long time before making a saccade.
@@ 3 Method
Our objective is to explore whether a bio-inspired model of visual learning can utilize the actual eye-tracking derived visual experience of toddlers to develop robust object representations. To mimic toddlers’ central visual experience, we use an eye-tracking dataset recorded during toddlers’ play sessions and extract parts of frames centered on the gaze location. For comparison, we also simulate different visual experiences following alternative gaze strategies. Then, we train a bio-inspired SSL model based on temporal slowness.
@@ 3.1 Toddler fixation dataset
The (Bambach et al., 2018) dataset contains head-camera videos recorded at 30 FPS and eye-tracking data for 38 dyads of toddlers/caregivers. All dyads play with the same set of 24 toys for 12 minutes on average. The children’s ages range from 12.3 to 24.3 months. For 30 dyads, a head-camera resolution of 480 $\times$ 640 pixels was used, while four dyads were recorded at 480 $\times$ 720 pixels and the remaining four at 240 $\times$ 320 pixels. The horizontal field of view covers 72 degrees. Figure 1A shows an example video frame with the gaze location (Bambach et al., 2018). In the following, we explain how we simulate different gaze strategies by deriving several datasets from these play sessions. Additionally, we include the anonymized information of all toddlers who participated in the study in Appendix C.
@@
Figure 1: Examples of visual sequences for each of our datasets. A. Raw frame from an egocentric video with the locations of our different croppings. Purple, orange, and green boxes representing gaze fixation, centroid fixation, and random fixation, respectively. The cross indicates the gaze location given by the eye-tracker. B-F. Example sequences for B- the Toddler fixation dataset; C- the Random fixation dataset; D- the Centroid fixation dataset; E the Objects fixation dataset and F- the Plain background dataset.
@@
Adult fixation dataset. We want to investigate the differences between gaze fixation in adults and toddlers and the consequences of these differences on learned representations. Thus, we also extract image patches around adults’ gaze fixation points following the procedure of the Toddler fixation dataset. Appendix A illustrates the gaze distribution of toddlers and adults.
Random fixation dataset. As a simple comparison dataset, we propose to simulate a completely random gaze strategy. We crop each frame at a uniform, random, and independently sampled location. Unlike Toddler/Adult fixation datasets, this dataset shows little spatio-temporal structure, and the cropped images are unlikely to contain well-centered objects. Figure 1D provides examples of random fixations.
Centroid fixation dataset. We also propose a stronger comparison dataset that simulates a static gaze within a moving head. Compared to the Random fixation dataset, this contains image patch sequences that are relatively stable over time. One possibility could be always to crop the center of the frames. However, we noticed that the head-camera was often misaligned with respect to the stationary position of the eyes, resulting in a mismatch between the center of the frames and the center of the camera wearer’s field of view. Thus, we rather use the centroid of the gaze fixation points (one for each video). To compute these centroids, we gather all gaze fixation points and calculate each video’s mean of their horizontal and vertical coordinates. Note that, despite the centroid positions being fixed, the continuous movement of the head changes the visible portions of the scene. Figure 1D presents a temporal sequence of the Centroid fixation dataset.
@@
Objects fixation dataset. This dataset was collected from the same video frames used in the Toddler fixation dataset. Images were automatically gaze-centered cropped with a 30-degree field of view, containing the target objects while minimizing the background interference (Tsutsui et al., 2021). This dataset contains 271,754 images. Figure 1E displays examples of images. The number of images per toy is depicted in Figure 2, which indicates that the dataset is imbalanced.
@@ 3.2 Self-supervised learning through time
To model the learning process of humans, we learn visual representations with a self-supervised model of temporal slowness, namely SimCLR-TT (Schneider et al., 2021). This algorithm is based on the state-of-the-art SimCLR method (Chen et al., 2020). SimCLR-TT samples an image $x_{t}$ at time $t$ and a temporally adjacent image $x_{t+\Delta{T}}$ at time $t+\Delta{T}$ and computes their respective embeddings $z_{t}$, $z_{t+\Delta{T}}$ with a deep neural network (e.g. ResNet). $\Delta{T}$ represents the time interval between two selected frames in SimCLR-TT; unless states otherwise, we set $\Delta{T}=\frac{1}{30}$ seconds. Then, SimCLR-TT minimizes
$\mathcal{L}\left(z_{t},z_{t+\Delta{T}}\right)=-\log\frac{\exp\left(\operatorname{sim}\left(z_{t},z_{t+\Delta{T}}\right)/\tau\right)}{\sum_{z_{k}\in\mathcal{B},k\neq t}\left[\exp\left(\operatorname{sim}\left(z_{t},z_{k}\right)/\tau\right)\right]},$ (1)
where $\mathcal{B}$ is a minibatch, ${\rm sim}(\cdot)$ is the cosine similarity and $\tau$ is the temperature hyper-parameter. Thus, SimCLR-TT maximizes the similarity between temporally adjacent representations (numerator) while keeping all representations dissimilar from each other (denominator). Figure 3 illustrates the learning process of SimCLR-TT.
Figure 3: Illustration of SimCLR-TT on the Toddler fixation dataset. Given that the videos operate at 30 FPS, the temporal gap $\Delta{T}=\frac{1}{30}$s.

### v2

# Toddlers’ Active Gaze Behavior Supports Self-Supervised Object Learning

@@ Abstract
Toddlers learn to recognize objects from different viewpoints with almost no supervision. Recent works argue that toddlers develop this ability by mapping close-in-time visual inputs to similar representations while interacting with objects. High acuity vision is only available in the central visual field, which may explain why toddlers (much like adults) constantly move around their gaze during such interactions. It is unclear whether/how much toddlers curate their visual experience through these eye movements to support their learning of object representations. In this work, we explore whether a bio-inspired visual learning model can harness toddlers’ gaze behavior during a play session to develop view-invariant object recognition. Exploiting head-mounted eye tracking during dyadic play, we simulate toddlers’ central visual field experience by cropping image regions centered on the gaze location. This visual stream feeds time-based self-supervised learning algorithms. Our experiments demonstrate that toddlers’ gaze strategy supports the learning of invariant object representations. Our analysis also reveals that the limited size of the central visual field where acuity is high is crucial for this. We further find that toddlers’ visual experience elicits more robust representations compared to adults’, mostly because toddlers look at objects they hold themselves for longer bouts. Overall, our work reveals how toddlers’ gaze behavior supports self-supervised learning of view-invariant object recognition.
@@ 1 Introduction
Toddlers learn visual representations that support the recognition of object instances observed from different viewpoints within their first year of life (Kraebel & Gerhardstein, 2006; Ayzenberg & Behrmann, 2024). This early emergence of view-invariant recognition and the ease with which adults perform this skill hide the complexities of learning it. Images reaching the retina vary drastically when objects are turned in depth. Even state-of-the-art machine learning methods still make absurd recognition mistakes when faced with unusual viewpoints of objects (Dong et al., 2022; Abbas & Deny, 2023; Ruan et al., 2023). This raises the question of what learning mechanisms support such view-invariant recognition in humans.
@@
One of the main theories posits that the development of view-invariant object recognition rests on the brain’s ability to construct visual representations that slowly change over time (Földiák, 1991; Li & DiCarlo, 2008; Miyashita, 1988). The main idea is that learners abundantly manipulate (or walk around) objects while watching them, giving access to different views of a single object over a short period of time. By learning slowly changing representations, a learner discards rapidly changing information from an image (here, information about the view) and naturally builds view-invariant representations. Following this idea, recent computational studies proposed to simulate humans’ visual experience by generating or curating large-scale temporal sequences of rotating objects (Aubret et al., 2022; Schneider et al., 2021; Yu et al., 2023); they confirm that learning slowly changing representations induces view-invariant object recognition. However, it is currently unclear if and how a toddler’s actual gaze behavior supports this learning mechanism.
@@
The significance of active gaze behavior stems from the limited area of high-acuity vision in humans. This area, known as the central visual field, covers only a few degrees of visual angle, but it dominates the extraction of semantic information in brain regions responsible for object recognition (Quaia & Krauzlis, 2024; Yu et al., 2015). However, such a small area of the visual field may be semantically unstable over time, as humans make three saccades per second on average. Then again, toddlers curate their own visual experience; compared to adults, objects held by toddlers appear bigger in the field of view due to their shorter arms (Bambach et al., 2018), select simpler stimuli (Anderson et al., 2024) and their visual inputs semantically change on a slower timescale (Sheybani et al., 2023). The latter point may be critical to make slowness-based learning operational.
@@
In this paper, we explore whether a bio-inspired model of visual learning can utilize the actual eye-tracking derived visual experience of toddlers to develop invariant object representations. For this, we leverage a dataset of head-camera recordings and gaze tracking from toddlers and adults during play sessions (Bambach et al., 2018). To simulate central visual experience, we crop image patches centered on tracked gaze locations. Then, we train previously introduced time-based self-supervised learning (SSL) models (Schneider et al., 2021). Our analysis shows that: a) toddlers’ gaze strategy boosts visual learning in comparison to several baselines; b) restricting learning to input from the central visual field improves object representations; and c) visual input from toddlers yields better representations than that from adults, which may be explained by toddlers looking longer at objects while manipulating them. In sum, our main contributions are:
  • We present the first ever study training SSL models on natural egocentric visual input derived from eye tracking in toddlers during play sessions.
  • We find that toddlers’ gaze strategy improves the learning of invariant object representations compared to several baselines.
  • We show that toddlers’ visual experience is more suitable for learning object representations through time-based SSL than adults’.
@@ 2 Related Work
Early computational studies found that slowness-based learning can extract representations of simple patterns that are invariant to position, size and rotation (Földiák, 1991; Wiskott & Sejnowski, 2002). Other works applied this principle to learn view-invariant object recognition (Wallis & Baddeley, 1997; Franzius et al., 2011; Einhäuser et al., 2005; Stringer et al., 2006). Recent advances in SSL allowed to scale the principle of temporal slowness to large sets of uncurated images of objects (Parthasarathy et al., 2022; Aubret et al., 2022; Schneider et al., 2021). This method was called SSL through time (SSLTT) (Aubret et al., 2022). On the machine learning side, SSLTT can boost category recognition (Aubret et al., 2024b; 2022; Sanyal et al., 2023), view-invariant object instance recognition (Schneider et al., 2021) and the alignment with human representations (Parthasarathy et al., 2023). On the cognitive modeling side, SSLTT can shape human-like inter-object semantic similarities (Aubret et al., 2024a) and combines well with visuo-language SSL to model object learning during dyadic play (Schaumlöffel et al., 2023). However, all these approaches use curated, synthetic, or third-person data, leaving unclear whether the statistical structure of toddlers’ actual visual experience, combined with temporal slowness, can indeed support object recognition. Another notable work studied the learning of view-invariant object representations in impoverished visual environments through the eyes of young chickens in a controlled rearing experiment (Pandey et al., 2024). In contrast, we apply SSLTT on natural visual inputs extracted from head cameras carried by toddlers and/or adults during play sessions.
@@
There is a recent surge in trained machine learning models on egocentric video datasets, including models of temporal slowness. For instance, the large-scale Ego4d dataset (Grauman et al., 2022) has been used for training vision models (Nair et al., 2022; Ma et al., ; Anderson et al., 2022). However, egocentric videos for toddlers have been missing (Anderson et al., 2022); this is a problem since existing research has found that the specific statistical structure of toddlers’ visual experience supports their learning (Sheybani et al., 2024; Bambach et al., 2017; Sheybani et al., 2023). The SAYcam dataset presents longitudinal recordings of 150 hours (on average) from each of the three participating children (Sullivan et al., 2021). With SAYcam, computational studies have shown that SSL methods can learn category recognition, with/without temporal slowness (Orhan et al., 2020; Orhan & Lake, 2024; Orhan et al., 2024). Another related work studies whether the temporal and developmental structure of toddlers’ visual experience supports category and action recognition, through temporal slowness (Sheybani et al., 2024). Yet, these computational studies neglect the gaze location and its associated behavioral strategy, as their datasets do not include the precise location of the individual’s gaze. We show in Section 4.2 that this is critical for learning good object representations.
@@
Our work extends previous approaches that also leverage the gaze location of a human to train vision models (Bambach et al., 2016; 2018). They also compare the quality of representations trained with toddlers’ versus adults’ experiences. However, these studies model the learning process with supervised learning, which is biologically implausible. This is important as, unlike bio-inspired self-supervised models that learn slowly changing representations, they are agnostic to the temporal structure of the visual experience, e.g., if toddlers look at an object for a long time before saccading to a different object.
@@ 3 Method
Our objective is to explore whether bio-inspired models of visual learning can utilize the actual eye-tracking derived visual experience of toddlers to develop robust object representations. To mimic toddlers’ central visual experience, we use an eye-tracking dataset recorded during toddlers’ play sessions and extract parts of frames centered on the gaze location. For comparison, we also simulate different visual experiences following alternative gaze strategies. Then, we train bio-inspired SSL models based on temporal slowness.
@@ 3.1 Toddler fixation dataset
The (Bambach et al., 2018) dataset contains head-camera videos recorded at 30 FPS and eye-tracking data for 38 dyads of toddlers/caregivers. All dyads play with the same set of 24 toys for 12 minutes on average. The children’s ages range from 12.3 to 24.3 months. For 30 dyads, a head-camera resolution of 640 $\times$ 480 pixels was used, while four dyads were recorded at 720 $\times$ 480 pixels and the remaining four at 320 $\times$ 240 pixels. The horizontal field of view covers 72 degrees. Figure 1A shows an example video frame with the gaze location (Bambach et al., 2018). In the following, we explain how we simulate different gaze strategies by deriving several datasets from these play sessions. Additionally, we include the anonymized information of all toddlers who participated in the study in Appendix C.
@@
Figure 1: Examples of visual sequences for each of our datasets. A. Raw frame from an egocentric video with the locations of our different croppings. Purple, orange, and green boxes representing gaze fixation, centroid fixation, and random fixation, respectively. The cross indicates the gaze location given by the eye-tracker. B-F. Example sequences for B- the Toddler fixation dataset; C- the Random fixation dataset; D- the Centroid fixation dataset; E the Objects fixation dataset and F- the Plain background dataset. Note that datasets E-F have been manually curated to only contain views of the target objects. This kind of oracle knowledge is not available to a naive learner.
@@
Adult fixation dataset. We want to investigate the differences between gaze fixation in adults and toddlers and the consequences of these differences on learned representations. Thus, we also extract image patches around adults’ gaze fixation points following the procedure of the Toddler fixation dataset. Appendix A illustrates the gaze distributions of toddlers and adults.
Random fixation dataset. As a simple comparison dataset, we propose to simulate a completely random gaze strategy. We crop each frame around a location that is sampled uniformly at random. Unlike the Toddler/Adult fixation datasets, this dataset shows little spatio-temporal structure, and the cropped images are unlikely to contain well-centered objects. Figure 1D provides example frames from the Random fixation dataset.
Centroid fixation dataset. We also propose a stronger comparison dataset that considers a human moving their head but not their eyes. This is an important comparison because it distills the effect of eye gaze. One possibility could be to always crop the center of the frames. However, we noticed that the head-camera was often misaligned with respect to the stationary position of the eyes, resulting in a mismatch between the center of the frames and the center of the camera wearer’s field of view (cf. Appendix A). Thus, we rather use the centroid of the gaze fixation points (one for each participant video). To compute these centroids, we gather all gaze fixation points and calculate each video’s mean of their horizontal and vertical coordinates. Note that, despite the centroid positions being fixed, the continuous movement of the head changes the visible portions of the scene. Nevertheless, compared to the Random fixation dataset, this set contains image patch sequences that are relatively stable over time. Figure 1D presents a temporal sequence of the Centroid fixation dataset.
@@
Objects fixation dataset. This dataset was collected from the same video frames used in the Toddler fixation dataset. Images were manually filtered such that toddlers looked at one of the target objects. From these frames crops with a 30-degree field of view around the gaze location were extracted, containing the target object while minimizing background interference (Bambach et al., 2018; Tsutsui et al., 2021). This dataset contains 271,754 images. Figure 1E displays examples of images. The number of images per toy is depicted in Figure 2, which indicates that the dataset is imbalanced. We conduct additional analysis on the class imbalance in Appendix B.3.
@@ 3.2 Self-supervised learning through time
To model the learning process of humans, we learn visual representations with a self-supervised model of temporal slowness, namely SimCLR-TT (Schneider et al., 2021). This algorithm is based on the state-of-the-art SimCLR method (Chen et al., 2020). SimCLR-TT samples an image $x_{t}$ at time $t$ and a temporally close image $x_{t+\Delta T}$ and computes their respective embeddings $z_{t}$, $z_{t+\Delta T}$ with a deep neural network (e.g. a ResNet). Unless stated otherwise, we set $\Delta{T}$ to the inverse of the camera’s frame rate, i.e., $\Delta{T}=\frac{1}{30}$ seconds. In Section 4.3 we show additional results varying $\Delta T$. Then, SimCLR-TT minimizes
$\color[rgb]{0,0,0}\mathcal{L}\left(z_{t},z_{t+\Delta T}\right)=-\log\frac{\exp\left(\operatorname{sim}\left(z_{t},z_{t+\Delta T}\right)/\tau\right)}{\sum_{z_{k}\in\mathcal{B},k\neq t}\left[\exp\left(\operatorname{sim}\left(z_{t},z_{k}\right)/\tau\right)\right]},$ (1)
where $\mathcal{B}$ is a minibatch, ${\rm sim}(\cdot)$ is the cosine similarity and $\tau$ is the temperature hyper-parameter. Here $k\neq t$ but $k=t+\Delta T$ is possible. Thus, SimCLR-TT maximizes the similarity between temporally close representations (numerator) while keeping all representations dissimilar from each other (denominator). Figure 3 illustrates the learning process of SimCLR-TT. In Appendix B.1 we also present results for BYOL-TT (Schneider et al., 2021).
Figure 3: Illustration of SimCLR-TT on the Toddler fixation dataset. By default the time interval $\Delta{T}=\frac{1}{30}$s corresponds to the inverse of the camera’s frame rate, but it can be increased to an integer multiple of this value.